OpenStudy (anonymous):

An airplane leaves an airport and flies due west 120 miles and then 150 miles in the direction S 39.17°W. How far is the plane from the airport (round to the nearest mile)?

OpenStudy (anonymous):



OpenStudy (anonymous):

cos B=(c^2+a^2-b^2)/2ca

OpenStudy (anonymous):

here angle B=90+39.17=129.17

OpenStudy (anonymous):

cos(90+39.17)=-sin39.17 c=120 b=? a=150 substitute the values and get the solution.